OpenStudy (anonymous):

-5/6x=-30 the answer is x =1 1/5 is this correct?

OpenStudy (anonymous):

i got 1/36

OpenStudy (anonymous):

i got 36=x

OpenStudy (anonymous):

i mean x=36 is the final answer, 100% sure

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

simply divide both sides of the equation by -5/6

OpenStudy (anonymous):

o I added and and got 35 the simplified to get 1 1/5

OpenStudy (anonymous):

if u do that^ u will be left with 1/x = 36

OpenStudy (anonymous):

i should of added 1 to 35. thats what happen.

OpenStudy (anonymous):

no you wont be left with 1/x=36

OpenStudy (anonymous):

x=(-30)/(-5/6) (-30/1)/(-5/6)=(-30/1)*(-6/5)=180/5=36

OpenStudy (anonymous):

o ok i understand now. thanks